OpenStudy (anonymous):

how to solve 8+4m=9m-7

OpenStudy (anonymous):

Try to move all the "m"s to one side, and all the numbers that are by themselves to the other. So, I'd add 7 to both sides first. Then I'd subtract 4m. Do these first and tell me what you get.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so like 7 - 8

OpenStudy (anonymous):

8+4m=9m-7 Add 7 to both sides. 8+4m + 7 = 9m-7 + 7 4m + 8 + 7 = 9m -7 + 7 4m + 15 = 9m + 0 4m + 15 = 9m Now subtract them by 4m.

OpenStudy (anonymous):

so 4m - 9m

OpenStudy (anonymous):

Actually no, you subtract 4m, so 9m - 4m

OpenStudy (anonymous):

So 4m + 15 -4m = 9m - 4m or 15 = 9m - 4m

OpenStudy (anonymous):

but dont you have to switch around like 9m - 4m = 15

OpenStudy (anonymous):

Doesn't matter. Equals means equals.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

So take it from there...

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

but before you do does it come up to being 9m divided by 9 =19 divided by 9

OpenStudy (anonymous):

No you don't need to divide by 9. You first want to combine the "m"s. So that's 9m - 4m = 5m. Make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In other words, your simplified equation is 15 = 5m

OpenStudy (anonymous):

Now you can divide both sides by 5

OpenStudy (anonymous):

and when you divide both sides by 5, you end up with 3=m (or m=3)

OpenStudy (anonymous):

what sides 15=5

OpenStudy (anonymous):

Sorry both sides of the equation. Taking a step back: 15 = 9m - 4m 15 = 5m \\[\frac{ 15 }{ 5 }=\frac{ 5m }{ 5 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 15 }{ 5 }=3\]

OpenStudy (anonymous):

\[\frac{ 5m }{ 5 }=m\]

OpenStudy (anonymous):

what about 19-13= -17p -5

OpenStudy (anonymous):

So this one isn't as complicated as it looks. You just have to cancel out all the numbers flying around this equation, then you'll end up with a negative P number that you'll have to turn into a positive.

OpenStudy (anonymous):

First step is to simplify it so that you're down to a single number equaling a multiple of P: On the left side, you have \[19-13=6\] So right away you can reduce the equation to \[6= -17p -5\]

OpenStudy (anonymous):

Next step is to get rid of the -5 on the right side, so you add 5 to both sides and get: \[11=-17p\]

OpenStudy (anonymous):

Now you have a much easier to read equation. So how do you solve for P? You have to divide both sides by -17, right? \[11\div -17=\frac{ -11 }{ 17 }\] and of course \[-17p \div-17=p\] so \[\frac{ -11 }{ 17 }=p\]